Zinc reacts with excess aqueous hydrochloric acid to produce hydrogen.

1. Write a balanced chemical equation for the reaction, using the smallest integer coefficients possible. Write states in brackets. (Hint: ZnCl2 is the second product.) Please include units. (aq, s, l)

2. Calculate the mass of pure Zn that will furnish 24.8 L of hydrogen at a pressure of 0.947 atm and a temperature of 32.7°C.

_______ g.
1
Expert's answer
2019-03-18T02:21:34-0400

1. Zn(s) + 2HCl(aq) → ZnCl2(aq) + H2(g)

2. n(H2) = PV/RT = (0.947atm*24.8L)/(305.85K*0.082atm*L/K*mol) = 0.936 mol

According to the equation n(Zn) = n(H2) = 0.936 mol

m(Zn) = n*M = 0.936mol*65g/mol = 60.84 g

60.84 g